OpenStudy (anonymous):

There were 300 buttons in a box. 140 were coloured red and there twice as many red buttons as green buttons. The remainder of the buttons were pink. what percentage of the buttons were pink?

OpenStudy (anonymous):

pink buttons =(300 - (140 +.5*140))300 *100 =30

OpenStudy (anonymous):

red 140 green = 140/2 = 70 pink = 300 - (140 + 70) = 90

OpenStudy (anonymous):

%age of pink buttons = 90/300 * 100% = 30%

OpenStudy (hoblos):

140 red 70 green 300-(140+70) = 90 pink percentage of pink = (90/300)*100 = 30%

OpenStudy (anonymous):

@JopHP yr red + green = 140 + 280 = 420!!!!!!

OpenStudy (anonymous):

Thanks JopHP ....☺

OpenStudy (anonymous):

sorry the calculator deceived me :(

OpenStudy (anonymous):

Don't worry, we all make mistakes sometimes....☺